OpenStudy (anonymous):

What is the equation for the line that passes through (1, -4) and is perpendicular to y = -2x +4?

OpenStudy (anonymous):

I get y = 2x-6, but supposedly the answer is y=0,5x-4,5. I'm thinking the book is wrong. Thoughts?

hartnn (hartnn):

what is the slope of given line y=-2x+4 ?? can u find slope ?

OpenStudy (anonymous):

well, according to the formula y = kx + m, it should be 2, given that it's perpendicular to -2

OpenStudy (anonymous):

the slope of y=-2x+4 is -2.. the slope of the perpendicular line is 1/2

OpenStudy (anonymous):

since m1*m2 = -1 given m1 is perpendicular to m2

OpenStudy (anonymous):

so now you have a slope and a point .. do you know how to find the line equation ?

OpenStudy (anonymous):

Hmm, sorry how did you calculate the slope?

OpenStudy (anonymous):

the slope of the given line is -2 right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now slopes of perpendicular lines satisfies : m1 * m2 = -1 now we know m1 = -2 -2 * m2 = -1 m2 = 1/2

OpenStudy (anonymous):

Totally confused. We only have m1, which is 4, where does m1 *m2 = -1 come from??

OpenStudy (anonymous):

m1 is -2 !! m - represents slope

OpenStudy (anonymous):

y=mx+n

OpenStudy (anonymous):

m2 - is the slope of the perpendicular line that we are looking for

OpenStudy (anonymous):

sorry if my letters\symbols are different from yours

OpenStudy (anonymous):

ah, right, yeah I would have called that k :) I'll try this again...

OpenStudy (anonymous):

ok now you have slope and a point so it should be easy for you

OpenStudy (anonymous):

gah, nope... Are you sure m1*m2= -1? If I draw (2/1) and (-2/1) they're perpendicular (unless I'm drawing them wrng), but 2 * (-2) = -4

OpenStudy (anonymous):

yes.. im sure http://www.purplemath.com/modules/slope2.htm

OpenStudy (anonymous):

lets do this together .. we found m2 = 1/2 now we have the point (1,-4) so the perpendicular line equation is : y2 = -4 + 0.5(x-1) y2 = -4 + 0.5x -0.5 = 0.5x -4.5x

OpenStudy (anonymous):

allright, I'm convinced :) thanks for you help time for a break for me I think

OpenStudy (anonymous):

yw